jumping from a height of 300 meters . At the time the cord lifted his brother back up , he was 23 meters above the ground . If Jamal started recording as soon as his brother fell , how much time elapsed when the cord snapped back ? Use f(t) = - 16t ^ 2 + c where is the height in meters
1
Expert's answer
2020-10-01T13:56:41-0400

"f(t)=-16t^2+300=23, t\geq0"


"16t^2=277"

"t=\pm\sqrt{\dfrac{277}{16}}"

Since "t\geq0," we take


"t=\sqrt{\dfrac{277}{16}}=\dfrac{\sqrt{277}}{4}\approx4.161(s)"